OpenStudy (anonymous):

Find the limit of the function algebraically lim x->-2 x^2-4/x+2

OpenStudy (sparrow2):

i think it's -4

OpenStudy (anonymous):

Could you elaborate?

OpenStudy (anonymous):

I forgot how to work these questions

OpenStudy (er.mohd.amir):

x2-4/(x+) =(x+2)*(x-2)/(x+2)= =x-2 take lim x-2=-2-2=-4

OpenStudy (anonymous):

Thank you!